# A Wearable Strain-Sensor-Based Shoulder Patch for Fatigue Detection in Bicep Curls

### v1

@@ III-C1 Experiment 1: Comparison of Changes in Strain Resistance Against sEMG
The result is presented in Fig. 10. The figure shows two $R_{norm}$ and sEMG RMS pair under two conditions: weight-free bicep curl and standard bicep curl. Our study observed that the maximum amplitude of the sEMG RMS from the BIC muscle is approximately 0.05 V during weight-free bicep curl, increasing to 0.15 V during standard bicep curl. In contrast, the corresponding changes in the resistance measured by the strain sensor were minimal.
@@
III-C2 Experiment 2: Comparison of Changes in Strain Resistance Against Shoulder Elevation Kinematics
This experiment aimed to determine whether the strain sensor could accurately capture the changes corresponding to those in the shoulder elevation trajectory. The overall experimental setup used was similar as illustrated in Fig. 8 with slight modifications on the sensor placement of the strain sensor on the AC joint as shown in Fig. 9B. The experiment records the subject’s shoulder kinematics and the strain sensor signal while conducting shoulder elevation with varying amplitudes.
The result is presented in Fig. 11. The figure shows $R_{norm}$ and the shoulder elevation trajectory during partial shoulder elevation and full shoulder elevation. Our results demonstrate that the resistance changes of our strain sensor correlate with shoulder joint movements. When the shoulder elevation ROM increased from 7 to 20 degrees, the maximum normalized resistance increased from 1.3 to 4.8, exhibiting a similar trend to the joint kinematics trajectory. The incremental ratio between the strain sensor signal and shoulder elevation ROM is approximately 1.4, indicating a proportional relationship between joint movement and skin deformation captured by the sensor.
@@ III-C2 Experiment 2: Comparison of Changes in Strain Resistance Against Shoulder Elevation Kinematics
III-C3 Experiment 3: Changes in Strain Resistance During Continuous Bicep Curls
@@ III-D1 Subjects
Subject Age (year) Height (cm) Weight (kg) Exercise duration per week (hour)
@@ III-E2 Shoulder Elevation Kinematics
The joint kinematics trajectory is analyzed to identify peaks and troughs. The cycle amplitude between each consecutive peak and trough was calculated. An increase of 150% over the cycle amplitude observed during standard bicep curls is used as a threshold. When the cycle amplitude for three consecutive trough-peak-trough sequences exceeds this threshold, the timestamp of the first peak in this sequence was recorded as the onset of fatigue $t_{k}$ (purple line), as indicated by joint kinematics.
@@ IV-A Obtaining Fatigue Timestamp
In our study, fatigue time was successfully captured based on the sEMG signal from 8 out of 13 subjects. The time differences between $t_{e}$ and $t_{s}$ were calculated. The average difference between these two metrics was 15.35 seconds, ranging from 1.86 seconds to 32.2 seconds, excluding data from subjects where fatigue was undetected. For the undetected subjects, the maximal differences in RMS computed between the reference and other windows was approximately 1.5.
@@ IV-A2 Kinematics
In our study, fatigue time was successfully captured based on shoulder kinematics from 12 out of 13 subjects. The time differences between $t_{k}$ and $t_{s}$ was calculated. The average time difference was 21.2 seconds, ranging from 2.69 seconds to 65.8 seconds.
@@ IV-B Subject Study: Real-Time Detection
The algorithm successfully captured the fatigue time for all subjects, and the resulting fatigue timestamps indicate a tendency for early detection. Specifically, the average absolute time difference difference from $t_{s}$ was 21.65 seconds, ranging from 2.1 seconds to 65.3 seconds.
@@ IV-C Subject Study: Post Hoc Detection
The efficacy of the proposed Post Hoc method for detecting fatigue was assessed using an algorithm that reliably captured fatigue time for all participants. The fatigue time was successfully captured for all subjects. Results indicated a tendency for late detection, with an average absolute time difference compared with $t_{s}$ of 9.54 seconds, ranging from 0.63 seconds to 19.4 seconds. 12 out of 13 subjects’ fatigue time was detected from the Pan Tompskin algorithm method, while only subject 6 inherited the fatigue time from the Real-Time method.
In contrast with the Real-Time method, which demonstrates higher accuracy for subjects with weekly exercise duration, the Post Hoc method does not prefer any specific exercise group. The nature of selecting the later time significantly reduces the high error cases for some subjects (subjects 4, 5, and 11). Despite increasing the error for some subjects (subjects 1, 2, 3, 8, and 9), most of the difference remains between 0.26 seconds and 5.64 seconds, with one outlier of 10.66 seconds.
@@
Subject Subject declared fatigue time $t_{s}$ Time differences with subject’s declaration
Kinematics $t_{s}-t_{k}$ sEMG $t_{s}-t_{e}$ Real-Time $t_{s}-t_{r}$ Post Hoc $t_{s}-t_{p}$
@@
    Avr 1: This is the average value of the absolute time differences, including the non-detected data, substituting the non-detected data with 0.
@@
    Avr 2: This is the average value of the absolute time differences excluding the non-detected data.
@@ V Conclusion
The findings demonstrate that this approach offers a cost-effective solution for identifying fatigue-induced muscle compensation during exercise.
The results indicate that kinematics-based detection successfully captured the majority of the subjects’ fatigue points. Based on this, we hypothesize that integrating kinematic sensors, such as IMU, could significantly enhance the accuracy and robustness of fatigue detection. IMUs would provide supplementary information on body dynamics, enriching the data obtained from traditional strain sensors and potentially improving the overall effectiveness of the wearable device in real-time fatigue assessment. Building on these promising results, our team is currently developing a more advanced wearable device that incorporates an IMU. This integration aims to further enhance the accuracy and reliability of fatigue detection by providing additional data on body dynamics.
In addition, our team is developing a complete wearable unit that integrates the sensor with a more user-friendly design, eliminating the need for an Arduino and long connector wires that currently limit movement and make the system impractical for home-based use.

### v2

@@ III-C2 Experiment 2: Comparison of Changes in Strain Resistance Against sEMG
The result is presented in Fig. 11. The figure shows two $R_{norm}$ and sEMG RMS pair under two conditions: weight-free bicep curl and standard bicep curl. Our study observed that the maximum amplitude of the sEMG RMS from the BIC muscle is approximately 0.05 V during weight-free bicep curl, increasing to 0.15 V during standard bicep curl. In contrast, the corresponding changes in the resistance measured by the strain sensor were minimal. Conducting Cross Correlation Analysis, the signal under the weight-free condition has a positive correlation of 0.58 with a lag of 1.43 seconds while the signals under the standard condition have a positive correlation of 0.49 with a lead of 0.3 seconds.
@@
III-C3 Experiment 3: Comparison of Changes in Strain Resistance Against Shoulder Elevation Kinematics
This experiment aimed to determine whether the strain sensor could accurately capture the changes corresponding to those in the shoulder elevation trajectory. The overall experimental setup used was similar to that in Fig. 8 with slight modifications on the sensor placement of the strain sensor on the AC joint as shown in Fig. 9B. The experiment records the subject’s shoulder kinematics and the strain sensor signal while conducting shoulder elevation with varying amplitudes.
The result is presented in Fig. 12. The figure shows $R_{norm}$ and the shoulder elevation trajectory during partial shoulder elevation and full shoulder elevation. Our results demonstrate that the resistance changes of our strain sensor correlate with shoulder joint movements. When the shoulder elevation ROM increased from 7 to 20 degrees, the maximum normalized resistance increased from 1.3 to 4.8, exhibiting a similar trend to the joint kinematics trajectory. The incremental ratio between the strain sensor signal and shoulder elevation ROM is approximately 1.4, indicating a proportional relationship between joint movement and skin deformation captured by the sensor. Conducting Cross Correlation Analysis, the signal in both stages achieved a high correlation of 0.92 and 0.95 with a lag of 0.01 seconds.
In summary, the strain sensor signal exhibited a moderate correlation with sEMG measurements (r = 0.49), indicating a reasonable association between muscle activation and strain during the standard bicep curl. In contrast, a strong correlation was observed between the strain signal and shoulder kinematics (r = 0.95), suggesting that the sensor is highly sensitive to compensatory joint movements. These findings support the use of the strain sensor as an effective surrogate for detecting fatigue-induced compensation, particularly through observable kinematic changes.
@@ III-C3 Experiment 3: Comparison of Changes in Strain Resistance Against Shoulder Elevation Kinematics
III-C4 Experiment 4: Changes in Strain Resistance During Continuous Bicep Curls
@@ III-D1 Subjects
Subject
Age
Height
Weight
Exercise duration
per week (hour)
@@ III-E2 Shoulder Elevation Kinematics
The joint kinematics trajectory is analyzed to identify peaks and troughs. The cycle amplitude between each consecutive peak and trough was calculated. An increase of 150% over the cycle amplitude observed during standard bicep curls is used as a threshold. When the cycle amplitude for three consecutive trough-peak-trough sequences exceeds this threshold, the timestamp of the first peak in this sequence is recorded as the onset of fatigue $t_{k}$ (purple line), as indicated by joint kinematics.
@@ IV-A Obtaining Fatigue Timestamp
In our study, fatigue time was successfully captured based on the sEMG signal from 8 out of 13 subjects. The time differences between $t_{e}$ and $t_{s}$ were calculated. The average difference between these two metrics was 15.35 seconds, with a standard deviation of 19.50 seconds, ranging from 1.86 seconds to 32.2 seconds, excluding data from subjects where fatigue was undetected. For the undetected subjects, the maximal differences in RMS computed between the reference and other windows was approximately 1.5.
@@ IV-A2 Kinematics
In our study, fatigue time was successfully captured based on shoulder kinematics from 12 out of 13 subjects. The time differences between $t_{k}$ and $t_{s}$ was calculated. The average time difference was 21.2 seconds, with a standard deviation of 27.48 seconds, ranging from 2.69 seconds to 65.8 seconds.
@@ IV-B Subject Study: Real-Time Detection
The algorithm successfully captured the fatigue time for all subjects, and the resulting fatigue timestamps indicate a tendency for early detection. Specifically, the average absolute time difference from $t_{s}$ was 21.65 seconds, with a standard deviation of 21.80 seconds, ranging from 2.1 seconds to 65.3 seconds.
@@ IV-C Subject Study: Post Hoc Detection
The efficacy of the proposed Post Hoc method for detecting fatigue was assessed using an algorithm that reliably captured fatigue time for all participants. The fatigue time was successfully captured for all subjects. Results indicated a tendency for late detection, with an average absolute time difference compared with $t_{s}$ of 9.54 seconds, with a standard deviation of 7.70 seconds, ranging from 0.63 seconds to 19.4 seconds. 12 out of 13 subjects’ fatigue time was detected from the Pan Tompskin algorithm method, while only subject 6 inherited the fatigue time from the Real-Time method.
In contrast with the Real-Time method, which demonstrates higher accuracy for subjects with weekly exercise duration, the Post Hoc method does not prefer any specific exercise group. The nature of selecting the latter time significantly reduces the high error cases for some subjects (subjects 4, 5, and 11). Despite increasing the error for some subjects (subjects 1, 2, 3, 8, and 9), most of the difference remains between 0.26 seconds and 5.64 seconds, with one outlier of 10.66 seconds.
@@
Subject
Subject
declared
fatigue
time $t_{s}$
Time differences
with subject’s declaration (s)
Kinematics
$t_{s}-t_{k}$
sEMG
$t_{s}-t_{e}$
Real-Time
$t_{s}-t_{r}$
Post Hoc
$t_{s}-t_{p}$
@@
    NA: Not detected.
@@
    Avr 1: This is the average value and the standard deviation of the absolute time differences, including the non-detected data, substituting the non-detected data with 0.
  • c
    Avr 2: This is the average value and the standard deviation of the absolute time differences, excluding the non-detected data.
  • d
    The values in Avr 1 and Avr 2 are reported as Mean (Standard Deviation).
@@ IV-D3 Exercise Frequency Factor
IV-E Home-based Application
The sensor system is designed to facilitate a direct transition from laboratory settings to in-home use, ensuring maximum user-friendliness. Setup requires minimal effort and can be performed with assistance from a caregiver or family member without medical knowledge. The AC joint can be easily located by identifying the small protrusion on the shoulder. Once identified, the sensor application involves simply placing a layer of kinesiology tape as a sacrifice layer over this area, positioning the sensor on top, and firmly securing the connection wire on the back of the patient, thereby rendering the sensor system ready for use.
@@ V Conclusion
The findings demonstrate that this approach offers a cost-effective solution for identifying fatigue-induced muscle compensation during exercise. The results indicate that kinematics-based detection successfully captured the majority of the subjects’ fatigue points. Based on this, we hypothesize that integrating kinematic sensors, such as IMU, could significantly enhance the accuracy and robustness of fatigue detection. IMUs would provide supplementary information on body dynamics, enriching the data obtained from traditional strain sensors and potentially improving the overall effectiveness of the wearable device in real-time fatigue assessment. Building on these promising results, our team is currently developing a more advanced wearable device that incorporates an IMU. This integration aims to further enhance the accuracy and reliability of fatigue detection by providing additional data on body dynamics.
In addition, our team is developing a complete wearable unit that integrates the sensor with a more user-friendly design, eliminating the need for an Arduino and long connector wires that currently limit movement and make the system impractical for home-based use. Future iterations will consider some commercial low-power microcontrollers, such as ARM Cortex M0, STM32L, and ESP32-S2, to enable a more compact and efficient solution for wearable applications.